OpenStudy (rainbow_dash):

A jar contains 3 chocolate cookies, 5 peanut butter cookies, and 6 coconut cookies. If 3 cookies are selected in succession, what is the probability of selecting chocolate, then peanut butter, and then coconut cookies, if replacement occurs each time?

OpenStudy (anonymous):

there are 14 cookies total right? so probability of getting a chocolate is \[\frac{3}{14}\] of getting a peanut butter is \[\frac{5}{14}\] and of getting a coconut is \[\frac{6}{14}\] multiply these together to get your answer

OpenStudy (rainbow_dash):

okay thank you :)